Solve each linear equation. Show your work and check you answer
9 - 2m/7 = 19
3(x - 0.87) - 2x = 4.98
Solve each problem
Fahrenheit temperature. Water boil at 212 F.
a) Use the accompanying graph to determine the Celsius temperature at which water boild
Celsius temperature 100

50
0
0 100 212
b) Find the Fahrenheir temperature of hot tap water at 70 C by solving the equation
70 = 5/9 (F - 32)

Answer by checkley71(8403)

9-2M/7=19 OR 2M/7=10 OR 2M=70 OR M=70/2 OR M=35
3(X-.87)-2X=4.98 OR 3X-2.61-2X=4.98 OR X=4.98+2.61 OR X=7.59
70=5/9(F-32) OR 70=(5/9)F-160/9 OR (5/9)F=70+160/9 OR (5/9)F=(630+160)/9 OR
45F=790*9 OR 45F=7110 OR F=158